OpenStudy (anonymous):

complete the square ? x2 + 8x = 11

OpenStudy (anonymous):

What would you add to both sides?

OpenStudy (anonymous):

idk , 8

OpenStudy (anonymous):

You take half of the x term so half of 8 = 4 Then you square that answer 4^2 = 16

OpenStudy (anonymous):

do you know what the x term refers to?

OpenStudy (anonymous):

no how many times do i have to say math is thee worse subject i dont know how to do this

OpenStudy (anonymous):

the x term refers to the number in front of the x

OpenStudy (anonymous):

you are asked to make the given equation into a square.

OpenStudy (anonymous):

x^2+8x-11 = 0

OpenStudy (anonymous):

do you know what the formula for (a+b)^2 is?

OpenStudy (anonymous):

in other words, \[(a+b)^{2} = ?\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay, \[(a+b)^{2} = a ^{2}+ 2ab +b ^{2}\]

OpenStudy (anonymous):

so the example that i just put up how would that work

OpenStudy (anonymous):

compare \[x ^{2}+8x - 11.\] and \[a ^{2}+2ab+b ^{2}\]

OpenStudy (anonymous):

x is a. 2ab is 8x

OpenStudy (anonymous):

so what must you add to -11 to make it into b^2?

OpenStudy (anonymous):

2ab

OpenStudy (anonymous):

okay try again. when you compare the two equations, you see that x = a. correct? and 2ab = 8x. so find the value of b.

OpenStudy (anonymous):

which grade are you in?

OpenStudy (anonymous):

idk , im srry im not Tht qood in math

OpenStudy (anonymous):

okay, 2ab = 8x, divide by 2a on both sides. => b = 8x/2a you know that x = a so what is the value of b?

OpenStudy (anonymous):

16 , i did 8*2 16

OpenStudy (anonymous):

i mean 4 ,

OpenStudy (anonymous):

okay good.

OpenStudy (anonymous):

thats riqqht

OpenStudy (anonymous):

so for (a+b)^2 to be equal to (x+4)^2, what must you add to x^2 +8x-11?

OpenStudy (anonymous):

expand (x+4)^2.

OpenStudy (anonymous):

maybe 2 0r 4 because i did 8\[\div\]2 qive me 4 then i square root that and qave me 2.

OpenStudy (anonymous):

what? you found out that x = a and 4 = b. so your complete square is (x+4)^2. so expand (x+4)^2 using the formula i gave earlier.

OpenStudy (anonymous):

okay \[(x+4)^{2}= x ^{2}+8x+16\]

OpenStudy (anonymous):

im in the 12th .

OpenStudy (anonymous):

you are in 12th grade and don't know what (a+b)^2 is ?That is very surprising. Anyway, compare the expanded version of (x+4)^2 i gave you with your original formula, x2+8x-11

OpenStudy (anonymous):

how much do you need to add to x2+8x-11 to get x2+8x+16?

OpenStudy (anonymous):

itss not because to mee math is hard for one and for two yu dont have to help mee someone else can if its like that im on here for help not for someone to tell im in the 12th i dont know what something is i guess maybe it dont get through to mee to many smart retricepeople wrong thing for me to come to flutter .

OpenStudy (anonymous):

im dont off this question anyway .

OpenStudy (anonymous):

make some sense. I did not want to offend you. I was just surprised. I thought basic algebra was taught in 8th grade or something.

OpenStudy (anonymous):

you know what flutter you and your negative attititude if yu aint going to help go some where with you ignorgant self and it probably was and wht its 2011 not 2007

OpenStudy (anonymous):

okay, have a pleasant day.

OpenStudy (anonymous):

you to , feather ! *

OpenStudy (anonymous):

Ah, well at least I'm not stupid.